### v1

@@ Abstract
Knowledge graphs are widely used in industrial applications, making error detection crucial for ensuring the reliability of downstream applications. Existing error detection methods often fail to effectively leverage fine-grained subgraph information and rely solely on fixed graph structures, while also lacking transparency in their decision-making processes, which results in suboptimal detection performance. In this paper, we propose a novel Multi-Agent framework for Knowledge Graph Error Detection (MAKGED) that utilizes multiple large language models (LLMs) in a collaborative setting. By concatenating fine-grained, bidirectional subgraph embeddings with LLM-based query embeddings during training, our framework integrates these representations to produce four specialized agents. These agents utilize subgraph information from different dimensions to engage in multi-round discussions, thereby improving error detection accuracy and ensuring a transparent decision-making process. Extensive experiments on FB15K and WN18RR demonstrate that MAKGED outperforms state-of-the-art methods, enhancing the accuracy and robustness of KG evaluation. For specific industrial scenarios, our framework can facilitate the training of specialized agents using domain-specific knowledge graphs for error detection, which highlights the potential industrial application value of our framework. Our code and datasets are available at https://github.com/kse-ElEvEn/MAKGED.
@@ 1 Introduction
Knowledge graphs (KGs) [1] represent facts in the real world as triples, such as (Paris, capital_of, France), facilitating the organization and scaling of information [5], has gained paramount importance in knowledge-based systems, such as retrieval-augmented generation and recommendation systems [20]. However, most large-scale KGs [1], built using rule-based methods to extract web data, often contain noisy or incorrect triples. For instance, the widely used knowledge graph NELL [15] contains around 600K incorrect triples, which account for 26% of the set of triples in NELL. Most knowledge graph-driven tasks assume all triples are correct [15], overlooking the impact of errors, which significantly degrades the performance of downstream tasks. This highlights the urgent need for effective KG error detection algorithms.
@@
Despite long-standing research interest, the challenge of developing effective KG error detection algorithms has not been well solved [9, 7, 13]. Existing approaches can be broadly categorized into two types based on their evidence utilization. First, explicit evidence-based methods extract paths or subgraphs from the knowledge graph that support the truth of triples and use these explicit evidences to predict the credibility of triples [16]. Second, embedding-based methods predict the credibility of triples by embedding entities and relations into vector space and calculating the embedded representation of paths or subgraphs [9, 7]. Moreover, when combined with contrastive learning and pre-trained models, their performance has shown significant improvement [24, 13, 14].
@@
However, these methods have two main limitations: (1) They evaluate triples from a single viewpoint, leading to incomplete evaluations and inaccurate confidence scores. In addition, these methods often rely on a fixed graph structure and specific textual representations for evaluation, meaning they use predefined, static connections between entities without dynamically adapting to the context of the triples being evaluated. As a result, they perform poorly when the graph structure or textual representations change. (2) They lack transparency in the evaluation process, typically providing a single confidence score without sufficient context about how decisions are made. These two limitations hinder error identification and impede model improvement.
@@
To address the limitations of existing methods, we introduce a multi-agent framework, as shown in Figure 1. For each triple in the KG, we assign two agents to the head and tail entities. The Forward Agent collects subgraphs with the entity as the head, and the Backward Agent collects subgraphs with it as the tail, integrating multiple perspectives for error detection. Then, we process the collected subgraphs using a Graph Convolutional Network (GCN) for structural features and a LLM for semantic features. By concatenating the GCN and LLM embeddings, we combine structural and semantic information, leveraging the LLM’s text generation capabilities for error detection.
@@
Moreover, to address the lack of a clear decision-making process during evaluation, we also introduce a structured discussion and voting mechanism. The multiple agents independently evaluate each triple and then engage in multiple rounds of discussion until they reach a final decision. This method enhances detection accuracy and increases transparency in the decision-making process. At the same time, our framework can leverage domain-specific knowledge graphs to train specialized agents for industrial scenarios, improving error detection accuracy. It addresses the limitations of existing LLMs, which struggle with domain-specific knowledge, thereby delivering significant industrial value. The main contribution are summarized as follows:
@@
    To our knowledge, we are the first to introduce a multi-agent framework, MAKGED, for knowledge graph error detection. Agents engage in three rounds of discussion and vote to determine triple correctness, enhancing the clarity and accuracy of decisions through collaborative evaluation.
@@
    We build datasets with simulated graph noise, and experiments show that MAKGED improves accuracy by 0.73% on FB15K and 6.62% on WN18RR compared to the state-of-the-art methods.
@@ 2 Related Work
Implicit Evidence : These methods use embedding models to map entities and relations into vector spaces for credibility evaluation. Examples include: CKRL [21] optimized triple credibility by considering local and global factors during embedding learning. KGTtm [9] combined random walks and embeddings, using MLP for scoring. Recent methods also use pre-trained language models and contrastive learning: KG-BERT [24] enriches semantic representations with pre-trained models. SeSICL [14], StAR [19], and CCA [13] align textual and structural representations via contrastive learning. These methods improve performance but rely on fixed graph structures, limiting adaptability to new or unseen structures.
@@ 3.3 Agent Construction
We construct four agents based on above subgraphs: Head_Forward_Agent, Head_Backward_Agent, Tail_Forward_Agent and Tail_Backward_Agent. Each agent analyzes the corresponding subgraph for the triple $(h,r,t)$, enabling a multi-angle evaluation of the triple by considering both head and tail entities’ forward and backward contexts.
@@ 4.1 Design of the Framework
In our framework, we employ multiple LLMs working collaboratively to detect errors in knowledge graphs. Using the structural information of the graph, we construct four bidirectional subgraph agents for both the head and tail entities. These agents analyze the contextual information of triples from different perspectives, and a final decision on the correctness of the triples is made through a voting mechanism. The detailed explanation of this process is provided below:
Bidirectional Subgraph Agents: In our framework, we design four bidirectional subgraph agents to evaluate triples in the knowledge graph. Each of these agents is responsible for analyzing triples from a specific directional perspective, including the Head_Forward_Agent and Head_Backward_Agent for the head entity, and the Tail_Forward_Agent and Tail_Backward_Agent for the tail entity, as illustrated in Figure 1.

### v2

@@ Abstract
Knowledge graphs are widely used in industrial applications, making error detection crucial for ensuring the reliability of downstream applications. Existing error detection methods often fail to effectively utilize fine-grained subgraph information and rely solely on fixed graph structures, while also lacking transparency in their decision-making processes, which results in suboptimal detection performance. In this paper, we propose a novel Multi-Agent framework for Knowledge Graph Error Detection (MAKGED) that utilizes multiple large language models (LLMs) in a collaborative setting. By concatenating fine-grained, bidirectional subgraph embeddings with LLM-based query embeddings during training, our framework integrates these representations to produce four specialized agents. These agents utilize subgraph information from different dimensions to engage in multi-round discussions, thereby improving error detection accuracy and ensuring a transparent decision-making process. Extensive experiments on FB15K and WN18RR demonstrate that MAKGED outperforms state-of-the-art methods, enhancing the accuracy and robustness of KG evaluation. For specific industrial scenarios, our framework can facilitate the training of specialized agents using domain-specific knowledge graphs for error detection, which highlights the potential industrial application value of our framework. Our code and datasets are available at https://github.com/kse-ElEvEn/MAKGED.
@@ 1 Introduction
Knowledge graphs (KGs) [1] represent facts in the real world as triples, such as (Paris, capital_of, France), facilitating the organization and scaling of information [5], and have gained paramount importance in knowledge-based systems, such as retrieval-augmented generation and recommendation systems [20]. However, most large-scale KGs [1], built using rule-based methods and statistics-based methods to extract web data, often contain noisy or incorrect triples. For instance, the widely used knowledge graph NELL [15] contains around 600K incorrect triples, which account for 26% of the set of triples in NELL. Most knowledge graph-driven tasks assume all triples are correct [15], overlooking the impact of errors, which significantly degrades the performance of downstream tasks. This highlights the urgent need for effective KG error detection methods.
@@
Existing methods can be broadly categorized into two types based on their evidence utilization. First, explicit evidence-based methods extract paths or subgraphs from the knowledge graph that support the truth of triples and use these explicit evidences to predict the credibility of triples [16]. Second, embedding-based methods predict the credibility of triples by embedding entities and relations into vector space and calculating the embedded representation of paths or subgraphs [9, 7]. Moreover, when combined with contrastive learning and pre-trained models, their performance has shown significant improvement [24, 13, 14].
@@
However, these methods have two main limitations: (1) They evaluate triples from a single, fixed perspective, either relying on static structural patterns (e.g., static graph embeddings) or a unidirectional semantic method (e.g., text embeddings). This means they assume predefined, unchanging connections between entities, without considering the dynamic context of the triples being evaluated. As a result, they struggle to adapt when the graph structure or textual representations change. (2) They lack transparency in the evaluation process, typically providing a single confidence score without sufficient context about how decisions are made. These two limitations hinder error identification and impede model improvement.
@@
To address the limitations of existing methods, we introduce a multi-agent framework, as shown in Figure 1. For each triple in the KG, we assign two agents to the head and tail entities. The Forward Agent collects subgraphs with the entity as the head, and the Backward Agent collects subgraphs with it as the tail, integrating multiple perspectives for error detection. Then, we process the collected subgraphs using a Graph Convolutional Network (GCN) for structural features and an LLM for semantic features. By concatenating the GCN and LLM embeddings, we combine structural and semantic information, leveraging the LLM’s text generation capabilities for error detection.
@@
Moreover, to address the lack of transparency during evaluation, we introduce a structured discussion and voting mechanism. The multiple agents independently evaluate each triple and then engage in multiple rounds of discussion until they reach a final decision. This method enhances detection accuracy and increases transparency in the decision-making process. At the same time, our framework can utilize domain-specific knowledge graphs to train specialized agents for industrial scenarios, improving error detection accuracy. It addresses the limitations of existing LLMs, which struggle with domain-specific knowledge, thereby delivering significant industrial value. The main contribution are summarized as follows:
@@
    To our knowledge, we are the first to introduce a multi-agent framework, MAKGED, for knowledge graph error detection. Agents engage in several rounds of discussion and vote to determine triple correctness, thus enhancing the clarity and accuracy of decisions through collaborative evaluation.
@@
    We build datasets with simulated graph noise, and experiments demonstrate that MAKGED improves accuracy by 0.73% on FB15K and 6.62% on WN18RR compared to the state-of-the-art methods.
@@ 2 Related Work
Implicit Evidence : These methods use embedding models to map entities and relations into vector spaces for credibility evaluation. CKRL [21] optimized triple credibility by considering local and global factors during embedding learning. KGTtm [9] combined random walks and embeddings, using MLP for scoring. Recent methods also use pre-trained language models and contrastive learning: KG-BERT [24] enriches semantic representations with pre-trained models. SeSICL [14], StAR [19], and CCA [13] align textual and structural representations via contrastive learning. These methods improve performance but rely on fixed graph structures, limiting adaptability to new or unseen structures.
@@ 3.3 Agent Construction
We construct four agents based on the above subgraphs: Head_Forward_Agent, Head_Backward_Agent, Tail_Forward_Agent and Tail_Backward_Agent. Each agent analyzes the corresponding subgraph for the triple $(h,r,t)$, enabling a multi-angle evaluation of the triple by considering both head and tail entities’ forward and backward contexts.
@@ 4.1 Design of the Framework
In our framework, we employ multiple LLM-based agents working collaboratively to detect errors in KGs. Using the structural information of the graph, we construct four bidirectional subgraph agents for both the head and tail entities. These agents analyze the contextual information of triples from different perspectives, and a final decision on the correctness of the triples is made through a voting mechanism. The detailed explanation of this process is provided below:
Bidirectional Subgraph Agents: In our MAKGED framework, we design four bidirectional subgraph agents to evaluate triples in the knowledge graph. Each of these agents is responsible for analyzing triples from a specific directional perspective, including the Head_Forward_Agent and Head_Backward_Agent for the head entity, and the Tail_Forward_Agent and Tail_Backward_Agent for the tail entity, as illustrated in Figure 1.